# A Curious Case of Remarkable Resilience to Gradient Attacks via Fully Convolutional and Differentiable Front End with a Skip Connection

### v1

@@ Abstract
We experimented with front-end enhanced neural models where a frozen backbone classifier was prepended by a differentiable and fully convolutional model with a skip connection. By training such composite models using a small learning rate for one epoch (or less), we obtained models that retained the accuracy of the backbone classifier while being unusually resistant to gradient attacks including APGD and FAB-T attacks from the AutoAttack package.
We provided evidence that this was due to gradient masking: Although the gradient masking phenomenon is not new, the degree of masking was quite remarkable for fully differentiable models that did not have gradient-shattering components such as JPEG compression or components that are expected to cause diminishing gradients. The training recipe to produce such models was remarkably stable and reproducible as well: We applied it to three datasets (CIFAR10, CIFAR100, and ImageNet) and several types of models (including recently proposed vision Transformers) without a single failure case.
Although black box attacks such as the SQUARE attack and the zero-order PGD can be partially effective against gradient masking, these attacks are easily defeated by combining gradient-masking models into simple randomized ensembles. We estimate that these ensembles achieve near-SOTA AutoAttack accuracy on CIFAR10, CIFAR100, and ImageNet (while retaining virtually all the clean accuracy of the original classifiers) despite having virtually zero accuracy under adaptive attacks. Quite interesting, adversarial training of the backbone classifier can further increase resistance of the front-end enhanced model to gradient attacks. On CIFAR10, the respective randomized ensemble achieved 90.8±$\pm 2.5$% (99% CI) accuracy under AutoAttack while having only 18.2±$\pm 3.6$% accuracy under the adaptive attack.
We do not aim to establish SOTA in adversarial robustness. Instead, our paper makes methodological contributions and further supports the thesis that adaptive attacks designed with the complete knowledge of model architecture are crucial in demonstrating model robustness and that even the so-called white-box gradient attacks can have limited applicability. Although gradient attacks can be complemented with black-box attack such as the SQUARE attack or the zero-order PGD, black-box attacks can be weak against randomized ensembles, e.g., when ensemble models mask gradients.
Code and instructions to reproduce key results is available. https://github.com/searchivarius/curious_case_of_gradient_masking/
**footnotetext: Work done outside of the scope of employment, primarily, while being employed by Bosch.
@@ 1 Introduction
In the field of adversarial machine learning, the AutoAttack package has gained quite a bit of popularity as a standard approach to evaluate model robustness. According to Google Scholar, by the end of 2023 core AutoAttack papers (Croce & Hein, 2020a; b; Andriushchenko et al., 2020) were cited about 2500 times. The AutoAttack is the “engine” of the RobustBench benchmark Croce et al. (2021), which is a popular benchmark tracking progress in adversarial machine learning. However, there has been little to no discussion in the literature about reliability of AutoAttack itself as well as its individual “components”. Tramèr et al. (2020) noted in passing that AutoAttack overestimated robust accuracy of several models, which had nearly zero accuracy under adaptive attacks, but they did not discuss the actual failure modes of the AutoAttack.
@@
To partially address this gap, we carry out a case study where the AutoAttack is applied to front-end “enhanced” models. We found such composite models as well as their randomized ensembles to be surprisingly resistant to standard projected gradient (PGD) attacks as well as to gradient attacks from the AutoAttack package (Croce & Hein, 2020a), which we attributed to gradient masking (and provided supporting evidence). This includes the FAB attack that was claimed to be resilient to gradient masking (Croce & Hein, 2020b). Despite the fact that RobustBench does not accept randomized models (Croce et al., 2021), users may apply AutoAttack to randomized defenses unaware that AutoAttack can fail against them (in fact, this was our experience). Moreover, the AutoAttack package was shown to be effective against complicated randomized defences (Croce & Hein, 2020a). Thus, it was reasonable to expect the AutoAttack to be successful against a simple randomized ensemble as well. When AutoAttack detects a randomized defense it only produces an easy-to-miss warning and suggests using the randomized version of the attack. In that, the randomized version of the AutoAttack has only gradient attacks, which—we find—can be ineffective in the presence of gradient masking.
@@
Although the gradient masking phenomenon is not new, the degree of this masking is quite remarkable for fully differentiable models that do not have gradient-shattering components such as JPEG compression or components that are expected to cause diminishing gradients (we also carried out sanity checks to ensure there were no numerical anomalies, see § A.1). In particular, the models retain full accuracy on original (non-adversarial) data, but a “basic” 50-step PGD with 5 restarts degrades performance of a front-end enhanced model by only a few percent, despite it is possible to circumvent gradient masking with nearly 100% success rate (using a variant of the backward pass differentiable approximation Athalye et al. 2018). Although black box attacks such as the SQUARE attack and the zero-order PGD can be effective against gradient masking, these attacks are easily defeated by combining gradient-masking models into simple randomized ensembles. We estimate that these ensembles achieve near-SOTA AutoAttack accuracy on CIFAR10, CIFAR100, and ImageNet despite having virtually zero accuracy under adaptive attacks. Quite interesting, adversarial training of the backbone classifier can further increase resistance of the front-end enhanced model to gradient attacks. On CIFAR10, the respective randomized ensemble achieved 90.8±$\pm 2.5$% (99% CI) accuracy under AutoAttack.
@@
We do not aim to establish SOTA in adversarial robustness. Instead, our paper makes methodological contributions and further supports the thesis that adaptive attacks designed with the complete knowledge of model architecture are crucial in demonstrating model robustness (Tramèr et al., 2020). We nevertheless believe that automatic evaluation is a useful baseline and we hope that our findings can help improve robustness and reliability of such evaluations. Code and instructions to reproduce key results is available. [^1] https://github.com/searchivarius/curious_case_of_gradient_masking/
@@ 2 Related Work
Our paper focuses on a specific sub-field concerned with test-time attacks and defenses for neural-network classifiers on image data, which dates back only to 2013 (Szegedy et al., 2014; Biggio et al., 2013). For a detailed overview we address the reader to recent surveys (Chakraborty et al., 2021; Akhtar et al., 2021). However, a broader field of adversarial machine learning has a longer history with earlier applications such as spam and intrusion detection (Dalvi et al., 2004; Lowd & Meek, 2005; Biggio & Roli, 2018).
@@
Originally, most practical classifiers were linear and despite successes in attacking them there was a belief that attacking complex and non-linear models would be difficult (Biggio & Roli, 2018). This belief was disproved by Biggio et al. (2013) and Szegedy et al. (2014). As Biggio & Roli (2018) noted, adversarial machine learning gained quite a bit of momentum after the publication by Szegedy et al. (2014). In addition to showing that deep learning networks were vulnerable to small adversarial perturbations, Szegedy et al. (2014) found that these perturbations were (to a large degree) transferable among models.
@@
Since then, the field has been growing exponentially, with more than eight thousand papers uploaded to arXiv from the beginning of 2018 to the end of 2023 (Carlini, 2019). Despite such an explosive growth, adversarial training is arguably the only reliable defense: which, nevertheless, does not provide theoretical guarantees: Some assurances can be provided by provable defenses, but they have limited applicability (Croce & Hein, 2020a; Akhtar et al., 2021).
@@
Quite a few additional empirical (sometimes sophisticated) defenses have been proposed, but they have been routinely broken using strong adaptive attacks (Athalye et al., 2018; Tramèr et al., 2020), which motivated development of automatic evaluation suits, which can reduce manual effort often required to execute adaptive attacks. Among them, AutoAttack (Croce & Hein, 2020b) is arguably the most popular package, which includes several parameter-free gradient attacks (including a newly proposed FAB (Croce & Hein, 2020b) as well as a query-efficient black-box SQUARE attack (Andriushchenko et al., 2020). About one year later Croce et al. (2021) created the RobustBench benchmark to track progress and adversarial machine learning where adversarial robustness is measured using the accuracy under AutoAttack.
@@
Despite Croce & Hein (2020a) advertised AutoAttack as a reliable attack, Tramèr et al. (2020) noted in passing that AutoAttack overestimated robust accuracy of models in several cases, but without providing details. In that, research on weaknesses of AutoAttack appears to be scarce. For example, Lorenz et al. (2021) question the general transferrability of RobustBench findings to more realistic settings, but not reliability of the attack itself. In particular, our finding regarding extreme gradient masking appears to be unknown. Methodologically, the most related paper is a work by Tramèr et al. (2020), who emphasize the importance of adaptive attacks. The main difference is that we focus on documenting one specific use of (extreme) gradient masking and simple randomization.
@@
Finally we note that gradient masking is typically associated with gradient shattering via non-differentiable components and randomization as well as components encouraging vanishing gradients (Athalye et al., 2018), but not with fully differentiable components with skip connections such the denoising front end used in our study (see see Figure 1 in § 4.1). We found only one case when a fully differentiable convolutional network became more resistant to PGD attacks as a result of using a special training procedure, namely, training with differential privacy (Boenisch et al., 2021; Tursynbek et al., 2020). However, the seemingly increased robustness was observed only on MNIST LeCun et al. (2010) and in the low accuracy zone (see Fig. 1a in the paper by Boenisch et al. 2021), whereas in our case gradient-masked models achieve accuracy of about 80-90% under a standard PGD attack while remaining full breakable (see Table 1).
@@ 3.1 Preliminaries
For training purposes the goodness of prediction is measured using a loss function $L$. The value of the loss function $L$ for input $x$ and true output $y$ is denoted as:
@@ 3.2 Adversarial Example and Threat Models
Adversarial examples are intentionally (and typically inconspicuously) modified inputs designed to fool a machine learning model. In the untargeted attack scenario, the adversary does not care about specific class outcomes as long as the model can be tricked to make an incorrect prediction. In the targeted attack scenario, the attacker objective is to make the model predict a specific class, e.g., to classify all images of pedestrians as motorcycles. The original unmodified inputs are often termed as clean data. The accuracy of the model on clean data is typically termed as clean accuracy. In contrast, the accuracy of the model on the adversarial examples is termed as robust accuracy.
In this paper, we consider adversarial input perturbations whose $L^{\infty}$ norm does not exceed a dataset-specific threshold $\varepsilon$: In other words, adversarial examples must be inside the ball ${\mathcal{B}}_{\varepsilon}(x)$, where $x$ is the original input. Different types of attacks require different levels of knowledge about the model and its test time behavior. The attacker can query the model (possibly limited number of times) by passing an input $x$ through the model and measuring one or more outcome such as:
@@
In some scenarios, the attacker can additionally know the model architecture and weights, which were produced during training. However, in the case of randomized models, it would not be possible to know test-time values of random parameters. In our study we consider three levels of access:
  • black-box attacks, which can obtain only the soft labels $h_{\theta}(x)$ and, consequently, the loss value $L(h_{\theta}(x),y)$;
  • gradient attacks, which can additionally access gradients $\nabla_{x}L(h_{\theta}(x),y)$; [^2] Note that these gradients are computed with respect to input image, rather than with respect to model parameters. Indeed, the objective of attack is to find the input maximizing the loss rather than the model parameters (that minimize the loss).

### v2

@@ Abstract
We experimented with front-end enhanced neural models where a differentiable and fully convolutional model with a skip connection is added before a frozen backbone classifier. By training such composite models using a small learning rate for about one epoch, we obtained models that retained the accuracy of the backbone classifier while being unusually resistant to gradient attacks—including APGD and FAB-T attacks from the AutoAttack package—which we attribute to gradient masking.
Although gradient masking is not new, the degree we observe is striking for fully differentiable models without obvious gradient-shattering—e.g., JPEG compression—or gradient-diminishing components. The training recipe to produce such models is also remarkably stable and reproducible: We applied it to three datasets (CIFAR10, CIFAR100, and ImageNet) and several modern architectures (including vision Transformers) without a single failure case.
While black-box attacks such as the SQUARE attack and zero-order PGD can partially overcome gradient masking, these attacks are easily defeated by simple randomized ensembles. We estimate that these ensembles achieve near-SOTA AutoAttack accuracy on CIFAR10, CIFAR100, and ImageNet (while retaining almost all clean accuracy of the original classifiers) despite having near-zero accuracy under adaptive attacks.
[…]
Code and instructions to reproduce key results are available. https://github.com/searchivarius/curious_case_of_gradient_masking
**footnotetext: Work done outside the scope of Amazon employment, primarily while employed by Bosch.****footnotetext: Work completed while employed by Bosch.
@@ 1 Introduction
In the field of adversarial machine learning, the AutoAttack package—henceforth simply AutoAttack—has gained considerable popularity as a standard approach to evaluate model robustness. According to Google Scholar, mid 2025 core AutoAttack papers (Croce & Hein, 2020a; b; Andriushchenko et al., 2020) had been cited about 4400 times. AutoAttack is the “engine” of the RobustBench benchmark Croce et al. (2021), which is a popular benchmark tracking progress in adversarial machine learning. However, the reliability of AutoAttack itself, as well as that of its individual “components,” has received little to no discussion in the literature. Tramèr et al. (2020) noted in passing that AutoAttack overestimated robust accuracy of several models, which had nearly zero accuracy under adaptive attacks, but they did not discuss the actual failure modes of AutoAttack.
To partially address this gap, we have carried out a case study where AutoAttack is applied to front-end “enhanced” models. We have found such composite models to be surprisingly resistant to standard projected gradient (PGD) attacks as well as to gradient attacks from AutoAttack (Croce & Hein, 2020a). We attribute this behavior to gradient masking (Athalye et al., 2018; Tramèr et al., 2020) and provide supporting evidence. Interestingly, the FAB attack from AutoAttack was previously shown to be resilient to gradient masking (Croce & Hein, 2020b). Yet, we have found it to be ineffective as well.
@@
Despite the fact that RobustBench does not accept randomized models (Croce et al., 2021), users may apply AutoAttack to randomized defenses unaware that AutoAttack can fail against them (in fact, this was our experience). Moreover, AutoAttack was shown to be effective against complex randomized defenses (Croce & Hein, 2020a). Thus, it was reasonable to expect AutoAttack to be successful against a simple randomized ensemble as well. When AutoAttack detects a randomized defense it only produces an easy-to-miss warning and suggests using the randomized version of the attack. However, the randomized version of AutoAttack has only gradient attacks, which—we find—can be ineffective in the presence of gradient masking.
@@
Even though the gradient masking phenomenon is not new, the degree of this masking is quite remarkable for fully differentiable models that do not have gradient-shattering components such as JPEG compression or components that are expected to cause diminishing gradients (we also carried out sanity checks to ensure there were no numerical anomalies, see § A.2). In particular, the models retain full accuracy on original (non-adversarial) data, but a “basic” 50-step PGD with 5 restarts degrades performance of a front-end enhanced model by only a few percent. At the same time it is possible to circumvent gradient masking with a nearly 100% success rate using a variant of the backward-pass differentiable approximation Athalye et al. 2018.
[…]
@@
We do not aim to establish SOTA in adversarial robustness. Our primary goal is to expose weaknesses of AutoAttack. In that, our results further support the argument that adaptive attacks designed with the complete knowledge of model architecture are crucial in demonstrating model robustness (Tramèr et al., 2020). We nevertheless believe that automatic evaluation is a useful baseline and we hope that our findings can help improve robustness and reliability of such evaluations. Beyond exposing weaknesses of AutoAttack, we also examine whether randomized ensembling can serve as a practical defense, and conclude the paper with a discussion of its potential. Code and instructions to reproduce key results are available. [^1] https://github.com/searchivarius/curious_case_of_gradient_masking
@@ 2 Related Work
Our paper focuses on a specific sub-field concerned with test-time attacks and defenses for neural-network classifiers on image data, which dates back only to 2013 (Szegedy et al., 2014; Biggio et al., 2013). For a detailed overview, we address the reader to recent surveys (Chakraborty et al., 2021; Akhtar et al., 2021). However, a broader field of adversarial machine learning has a longer history with earlier applications such as spam and intrusion detection (Dalvi et al., 2004; Lowd & Meek, 2005; Biggio & Roli, 2018).
@@
Originally, most practical classifiers were linear and despite successes in attacking them there was a belief that attacking complex and non-linear models would be difficult (Biggio & Roli, 2018). This assumption was partly rooted in the intuition that increased model complexity could make decision boundaries harder to approximate or exploit. However, neural networks—while non-linear—turned out to be vulnerable in ways similar to or even more so than their linear counterparts (Biggio et al., 2013; Szegedy et al., 2014). As Biggio & Roli (2018) noted, adversarial machine learning gained quite a bit of momentum after the publication by Szegedy et al. (2014). In addition to showing that deep neural networks were vulnerable to small adversarial perturbations, Szegedy et al. (2014) found that these perturbations were (to a large extent) transferable among models.
@@
Since then, the field has been growing exponentially, with more than twelve thousand papers uploaded to arXiv from the beginning of 2018 to mid-2025 (Carlini, 2019). Despite such explosive growth, adversarial training is arguably the only reliable defense. Nevertheless, it does not provide theoretical guarantees. Some assurances can be provided by provable defenses, but they have limited applicability (Croce & Hein, 2020a; Akhtar et al., 2021).
@@
Quite a few additional empirical (sometimes sophisticated) defenses have been proposed, but they have been routinely broken using strong adaptive attacks (Athalye et al., 2018; Tramèr et al., 2020). This, in turn, motivated development of automatic evaluation suites, which can reduce manual effort often required to execute adaptive attacks. Among them, AutoAttack (Croce & Hein, 2020b) is arguably the most popular package, which includes several hyperparameter-free gradient attacks as well as a query-efficient black-box SQUARE attack (Andriushchenko et al., 2020). About one year later, Croce et al. (2021) created the RobustBench benchmark to track progress in adversarial machine learning where adversarial robustness is measured using the accuracy under AutoAttack.
@@
Although Croce & Hein (2020a) advertised AutoAttack as a reliable attack, Tramèr et al. (2020) noted in passing that AutoAttack overestimated robust accuracy of models in several cases, but without providing details. In that, research on weaknesses of AutoAttack appears to be scarce. For example, Lorenz et al. (2021) question the general transferability of RobustBench findings to more realistic settings, but not reliability of the attack itself. In particular, our finding regarding extreme gradient masking appears to be unknown. Methodologically, the most related paper is by Tramèr et al. (2020), who emphasized the importance of adaptive attacks. The main difference is that we focus on documenting one specific use of (extreme) gradient masking and simple randomization.
@@
Note that gradient masking is typically associated with gradient shattering via non-differentiable components and randomization as well as components encouraging vanishing gradients (Athalye et al., 2018), but not with fully differentiable components with skip connections such as the denoising front end used in our study (see Figure 1 in § 4.1). We found only one case in which a fully differentiable convolutional network became more resistant to PGD attacks due to a special training procedure, namely, training with differential privacy (Boenisch et al., 2021; Tursynbek et al., 2020). However, the seemingly increased robustness was observed only on MNIST (LeCun et al., 2010) and in the low accuracy zone (see Fig. 1a in the paper by Boenisch et al. 2021). In contrast, in our case gradient-masked models achieve an accuracy of about 80-90% under a standard PGD attack while remaining fully breakable (see Table 1).
Randomized defenses fall under the broader category of moving-target defenses (Evans et al., 2011). Although our paper focuses on the weaknesses of AutoAttack rather than on proposing a new defense through randomization or obfuscation, it is worth discussing several typical approaches. This discussion covers some key approaches, but it is not intended to be exhaustive. Randomization and obfuscation broadly aim to prevent attackers from reproducing the exact data or model conditions used at prediction time. Existing approaches fall into two broad categories: the first involves random or pseudo-random modifications to the input data, while the second randomizes the model architecture, its parameters, or a combination thereof.
For example, Taran et al. (2019) use an ensemble of models, each operating on differently perturbed input data. A perturbation is created by (1) first transforming input using an invertible parameterized transformation $W_{i}$, (2) modifying the result using another parameterized transform $P_{k}$, and (3) then mapping it back using the inverse transformation $W_{i}^{-1}$. The choice of transformations is determined by secret keys unknown to the attacker. However, there is no inference-time randomization and the defense can potentially be circumvented by black-box attacks such as the SQUARE attack (Andriushchenko et al., 2020). In contrast, Qin et al. (2021) proposed to add Gaussian noise before passing inputs to the model—a method that aims to circumvent black-box attacks. Randomized smoothing also relies on randomly perturbing input data multiple times. It uses majority voting over noisy predictions to create a robust classifier, which also provides certification (Cohen et al., 2019). [^2] Certification is a formal guarantee that the classifier’s prediction does not change within a certain neighborhood, defined by a certification radius $R$. In other words, it ensures that the prediction is unaffected by any adversarial perturbation with radius at most $R$.
Model randomization includes a variety of defense strategies, with randomized ensembles being one of the simplest and most natural. In this approach, a subset of models (possibly just a single one) is selected at inference time to make a prediction. Even if the attacker has access to all models in the ensemble, they cannot know in advance which subset will be used for a given input. While we do not know what the first paper is that formally introduced randomized ensembles as a defense, Sengupta et al. (2019) proposed a game-theoretic extension of this idea, where optimal selection probabilities are computed based on each model’s performance on benign and adversarial inputs. A key limitation of their analysis is the assumption that the adversary generates attacks using only one model at a time. In practice, it may be possible to construct adversarial examples that transfer across all models—an attack strategy not considered in their work. For example, one can generate adversarial samples by attacking a model ensemble (Liu et al., 2017; Gubri et al., 2022).
That said, it is possible to reduce transferability among ensemble models. For example, Bui et al. (2021) proposed a method that combines minimizing the cross-entropy loss for one model in the ensemble on an adversarial example with maximizing the loss for other models that misclassify the same input. Importantly, it is not always the same model whose loss is maximized—each adversarial example is associated with a different weighting of loss components, depending on whether each model predicts it correctly. Although the additional maximization term discourages correct predictions, it also encourages mispredicting models to produce highly uncertain outputs, with post-softmax probabilities approaching a uniform distribution. As a result, the influence of such mispredictions on the ensemble’s overall prediction remains limited (the prediction is based on summing logits of all models in the ensemble).
Sharif et al. (2019) proposed a variant of adversarial training in which adversarial examples are generated using targeted attacks. Instead of searching for arbitrary perturbations that maximize the cross-entropy loss for the correct label $y$, they define a random label mapping m(y)m(y)—called a derangement—such that $m(y)\neq y$ for all $y$. In other words, $m$ is a randomly selected permutation that never maps a label to itself. Given an input $x$ with label $y$, they then construct adversarial examples that minimize the cross-entropy loss with respect to the target label $m(y)$, which is different from true label $y$. To promote diversity within the ensemble, a different derangement is sampled for each model.
Although not suggested for reducing intra-model transferability, Sitawarin et al. (2024) proposed light adversarial training where a target model is trained using adversarial examples generated for a set of publicly available models. This can also be applied to a multi-model scenario. To this end, at each training step one can sample a pair of models $A$ and $B$. Then, they generate an attack using the model $A$ and apply it to the adversarial training of model $B$.
Bai et al. (2024) observed that attacks exhibit limited transferability between an adversarially trained model and a model trained on original (clean) data—a finding we also confirm in our study (see § A.1). They exploited this insight by linearly combining the logits of the robust and standard classifiers (without retraining) and then applying a non-differentiable nonlinear transformation (with three parameters) to the result. The parameters of this transformation are optimized to maximize robustness.
A disadvantage of random ensembling is that it requires training and hosting of multiple models. A more efficient alternative is self-ensembling, which introduces randomness directly into a single model during training and/or inference. A basic example is test-time dropout, which—as we found—despite its simplicity can fool the SQUARE attack. It is also worth noting that randomization of input data with subsequent application of the same model—e.g., as in the randomized smoothing (Cohen et al., 2019)—can be seen as a form of self-ensembling as well.
A number of more sophisticated self-ensembling techniques have also been proposed. Liu et al. (2018) add a random noise layer before each convolutional layer, injecting Gaussian perturbations into the layer’s input. Dong et al. (2022) found that using different normalization approaches (e.g., batch normalization vs. instance normalization) reduces transferability of attacks, and they propose a random normalization aggregation approach, which randomly selects the normalization method for each layer that uses normalization.
Fu et al. (2021) proposed to train models using randomly selected precision. They combine random-precision inference (RPI) and random-precision training (RPT): RPI selects inference precision at run-time to exploit low attack transferability across precisions, while RPT randomly switches training precisions with switchable batch normalization to further reduce attack transferability and enhance robustness.
@@ 3.1 Preliminaries
For training purposes the accuracy of prediction is measured using a loss function $L$. The value of the loss function $L$ for input $x$ and true output $y$ is denoted as:
@@ 3.2 Adversarial Example and Threat Models
Adversarial examples are intentionally (and typically inconspicuously) modified inputs designed to fool a machine learning model. In the untargeted attack scenario, the adversary does not care about specific class outcomes as long as the model can be tricked to make an incorrect prediction. In the targeted attack scenario, an attacker’s objective is to make the model predict a specific class, e.g., to classify all images of pedestrians as motorcycles. The original unmodified inputs are often termed as clean data. The accuracy of the model on clean data is typically termed as clean accuracy. In contrast, the accuracy of the model on the adversarial examples is termed as robust accuracy.
In this paper, we consider adversarial input perturbations whose $L^{\infty}$ norm does not exceed a dataset-specific threshold $\varepsilon$. In other words, adversarial examples must be inside the ball ${\mathcal{B}}_{\varepsilon}(x)$, where $x$ is the original input. Different types of attacks require different levels of knowledge about the model and its test time behavior. The attacker can query the model (possibly a limited number of times) by passing an input $x$ through the model and measuring one or more outcomes such as:
@@
In some scenarios, the attacker can additionally know the model architecture and weights. However, in the case of randomized models, it would not be possible to know test-time values of random parameters. In our study, we consider three levels of access:
  • black-box attacks, which can obtain model predictions $\operatorname*{arg\,max}_{i}h_{\theta}(x)_{i}$ or the full soft label distribution $h_{\theta}(x)$. In the latter case, the attacker can also compute the loss value $L(h_{\theta}(x),y)$. Black-box attacks are commonly referred to as query-based attacks.;
  • gradient attacks, which can additionally access gradients $\nabla_{x}L(h_{\theta}(x),y)$; [^3] Note that these gradients are computed with respect to an input image, rather than with respect to model parameters. Indeed, the objective of the attack is to find the input maximizing the loss rather than the model parameters (that minimize the loss).